Problem:
find the constant of variation k
m varies directly with n; m = 144 when n = 8


Answer by ankor@dixie-net.com(12695)

m varies directly with n; m = 144 when n = 8
:
m = n*k
:
Substitute for m and n and find k:
144 = 8k
:
k = 144/8
:
k = 18
:
:
Check:
m = 18n
m = 18 * 8
m = 144